OpenStudy (turingtest):

I just got 71% on my first probability test :( Here is one I missed. The answers are not released yet so 1) I cannot confirm if we have correctly solved it and. 2) Do no look at this thread if you are still talking the test. To get a complete collection of stamps, Ann needs to have 5 specific (and different) kinds of stamps, denoted by \(k_1,…,k_5\). She starts out with no stamps. She then receives 5 stamps randomly from a lottery. Each of these 5 stamps has probability 1/5 of being of any particular kind \(k_i\), and the kinds of the different stamps she receives are independent.

OpenStudy (cakt07):

Lol im doing probabilty to, and i SUCK :P

OpenStudy (cakt07):

Sorry, i dont know this either, i really SUCK at probabilty, but when im done learning i can try my best to help

OpenStudy (turingtest):

Questions: a) What is the probability that Ann has a complete collection after receiving the 5 stamps from the lottery? (this one I got correctly as 24/625) b) Given that Ann does not have a complete collection after receiving the 5 stamps from the lottery, what is the probability that she is missing exactly one stamp? for this I tried 96/601 and (if i remember correctly) 24/601. neither were correct. Any thoughts?

OpenStudy (dulljackel09):

I just wanted to give you a medal bye

OpenStudy (turingtest):

yay, Professor Zarkon has arrived

OpenStudy (cakt07):

Professor Zarkon?

OpenStudy (cakt07):

Hola Professor!

OpenStudy (cakt07):

Um.... No, T_T Im goin with professor.

OpenStudy (cakt07):

At least 71% is a C in MY grade, If you in high-school, i think it might be either C- or D

OpenStudy (turingtest):

it's from MIT, so I think they give A's at 80%+, and give some room for bonus points, but still...

OpenStudy (anonymous):

i'm doing probability too though i suck horribly at it.

OpenStudy (anonymous):

aaaaa; 1*1 aaaab, c, d, e; 4*5 aaabc, bd, be, ce, de; 5*10 aabcd, bce, cde; 3*10 abcde; 1*5 530 different outcomes of (1/5)^5 each maybe? and 5 ways to pull out abcde? im sure somethings off there

OpenStudy (anonymous):

5^5 different positional ways ....

OpenStudy (anonymous):

and there are what, 5! ways to order S5 ?

OpenStudy (turingtest):

I may as well show my reasoning. for the first: \[P(All~~Stamps)=1\cdot\frac45\cdot\frac35\cdot\frac25\cdot\frac15=24/625\]for the second\[P(All~~Stamps-1|Not~~All~~Stamps)\\={P(All~~Stamps-1\cap Not~~All~~Stamps)\over P(Not~~All~~Stamps)}\\={1\cdot\frac45\cdot\frac35\cdot\frac25\cdot\frac45\over1-P(All~~Stamps)}={{4*4!\over5^4}\over\frac{601}{625}}=...\frac{96}{601}~~~?\]wait a minute... was my only error a calculator error? did I get the right answer just now?

OpenStudy (turingtest):

I may as well show my reasoning. for the first: \[P(All~~Stamps)=1\cdot\frac45\cdot\frac35\cdot\frac25\cdot\frac15=24/625\]for the second\[P(All~~Stamps-1|Not~~All~~Stamps)\\={P(All~~Stamps-1\cap Not~~All~~Stamps)\over P(Not~~All~~Stamps)}\\={1\cdot\frac45\cdot\frac35\cdot\frac25\cdot\frac45\over1-P(All~~Stamps)}={{4*4!\over5^4}\over\frac{601}{625}}=\frac{96}{601}\]

OpenStudy (turingtest):

pretty sure my problem is in the numerator...

OpenStudy (anonymous):

3.84 seems right to me for the first: 120/3125

OpenStudy (anonymous):

that looks correct

OpenStudy (anonymous):

Brass Monkey

OpenStudy (zarkon):

go into more detail on how you worked out your numerator

OpenStudy (turingtest):

i thought 1/1 chance for the first being a new coupon, 4/5 the next is new, 3/5 the next is new, 2/5 the next is new, then the last must be one of the prior 4, so 4/5 chance for the last one to already be in the set.... I don't think that covers the intersection of the sets, though

OpenStudy (zarkon):

think of it in terms of the multinomial distribution suppose you got 2 of the first kind and 0 of the last kind then the prob is \[\frac{5!}{2!1!1!1!0!}\left(\frac{1}{5}\right)^2\left(\frac{1}{5}\right)^1\left(\frac{1}{5}\right)^1\left(\frac{1}{5}\right)^1\left(\frac{1}{5}\right)^0\] now count the number of ways you can move the 2 and 0

OpenStudy (turingtest):

would that be\(\binom 5 2\) ways to move the 2 and 0? no they are distinct... I really didn't think of using multinomials

OpenStudy (zarkon):

no...the order of the 2 and 0 matters

OpenStudy (turingtest):

yeah I see that, one sec, blanking out lol

OpenStudy (turingtest):

the number of permutations of 2,1,1,1,0 are... ? that's what you are asking, right?

OpenStudy (zarkon):

yes

OpenStudy (turingtest):

why can I not think of a formula for this... sorry let me wallow in my stupidity for a sec

OpenStudy (zarkon):

how many choices do you have for the 2...then how many do you have for the 0

OpenStudy (turingtest):

5 and 5 respectively, and 5 choose 3 for the remaining 1's ?

OpenStudy (turingtest):

so we multiply those factors?

OpenStudy (zarkon):

just move the 2 and 0...the 1's then only have 1 way to be placed

OpenStudy (turingtest):

no 3! for the remaining 1's... I think

OpenStudy (turingtest):

oh, so just 25?

OpenStudy (zarkon):

;)...no...little smaller than that

OpenStudy (turingtest):

oh darn 20

OpenStudy (zarkon):

yes...so take 20 times the probability I wrote above

OpenStudy (turingtest):

5 places to put the 0, 4 to put the 2

OpenStudy (turingtest):

gotchya, never would have occurred to me to use multinomials. For the sake of completeness, do you see an outright error in my argument?

OpenStudy (zarkon):

yes...there are a few things wrong...the obvious one is that they you wrote it the 2nd choice cant be a repeat of the 1st (you have 4/5)...but it is possible that if the first two choices you get the same thing

OpenStudy (zarkon):

I hope you can translate that it didn't come out very well ;)

OpenStudy (turingtest):

but given she only missed one, doesn't the ordering of that not matter? if she gets stamp1 twice in a row, that's 1*1/5*4/5*3/5*2/5 ways, same thing, no? or does the condition change that? or am I making no sense at all?

OpenStudy (turingtest):

oh wait it's not the same lol

OpenStudy (zarkon):

for the first one you have 5 out of 5 possibilities...for the second you still have 5 out of 5 possibilities for the 3...now it gets hard since we would have to do cases...did you already have a repeat or not....

OpenStudy (turingtest):

I didn't mean ways, I meant chance above

OpenStudy (dulljackel09):

Its a nerd convention lol

OpenStudy (turingtest):

ok, I think I get the picture, I'll let y'all know theofficil solution when it is released

OpenStudy (turingtest):

I have a feeling the solution they give will not use multinomials, but we shall see :) Thanks, Doc

OpenStudy (zarkon):

I get...after simplifying ...\[\frac{240}{601}\approx.39993\]

OpenStudy (turingtest):

that seems like a high probability to me intuitively, that it's a 40% chance she's missing only one given she doesn't have them all, but those kinds of things are hard to estimate sometimes

OpenStudy (zarkon):

when I first did it, I didn't use the multimomial. I just thought it would be easier to explain the solution using them

OpenStudy (turingtest):

Thank you, your solution makes perfect sense. If you feel like typing out your other solution I'll be happy to see it, but if you don't feel like going to the trouble, no worries. Or if it uses anything past basic pmf introduction, which is where I am, I won't follow anyway.

OpenStudy (zarkon):

I don't think it would make much sense. ;)

OpenStudy (turingtest):

I trust that :D

OpenStudy (zarkon):

if you have matlab...or one of its clones you could try running this ------------------------- function p=stamp(n) t=0; w=0; for k=1:n x=ceil(5*rand(1,5)); y=sum(ismember(1:5,x)); if y<5 w=w+1; if y==4 t=t+1; end end end p=t/w; ------------------ it will give the conditional probability

OpenStudy (zarkon):

I like to check my answers by doing simulations

OpenStudy (turingtest):

Nice, unfortunately my copy of matlab suddenly crashed permanently and no longer compiles anything. Thanks for the CS approach though :)

OpenStudy (ybarrap):

Here's an online Matlab that is open to everyone - http://www.compileonline.com/execute_matlab_online.php There is also Octave, that is a free and open source version of Matlab - https://www.gnu.org/software/octave/ (doesn't have all the libraries of Matlab, but is good for home-grown code)

OpenStudy (ybarrap):

Nice work @zarkon!

OpenStudy (turingtest):

Here is the solution as provided by MIT, for all those who want to see another solution. Variety is, after all, the spice of life. Given that Ann does not start out with a complete collection, the number of unique sequences of Ann's stamps is \(5^5−5!\), each of which is still equally likely in the conditional universe. Thus, we need only count the number of sequences with exactly 1 missing stamp and then divide by \(5^5−5!\). There are 5 possibilities for the missing stamp, 4 possibilities for the duplicate stamp, \(\binom52\) ways to place the 2 duplicate stamps in the sequence of 5 stamps, and \(3!\) ways to place the remaining 3 stamps in the remaining 3 positions in the sequence. Therefore, the probability that Ann is missing one stamp is \[{5⋅4⋅\binom52⋅3!\over5^5−5!}={5!\binom52\over5^5−5!}=\frac{240}{601}≈0.39933. \]

OpenStudy (turingtest):

@Zarkon I like your explanation better